OpenStudy (christos):

Find amplitude and the phase shift of the function f(x) = -3sin (2x+Π/2) ANSWER: |A|=3;phase=-Π/4 I need to find how to solve that in order to find this answer

hero (hero):

Did you figure out the correct answer to the previous question?

OpenStudy (christos):

Yea, I still cant find this one though

hero (hero):

You didn't let me know what it was. How do I know you got it correct unless you tell me?

OpenStudy (christos):

It was Π/2

hero (hero):

Okay good.

hero (hero):

The amplitude of a sinusoidal function is the absolute value of A so: A = -3 so |A| = |-3| = 3 To find the Phase Shift of Bx + C, factor out B: 2x+Π/2 = 2(x + Π/4) Thus the Phase Shift is Π/4

hero (hero):

Basically, Bx + C = B(x + C/B). It's a little tricky to understand but you're smart.

hero (hero):

Did you understand any of that?

OpenStudy (christos):

But the answer to the solution in the book is written as -Π/4 . I understood everything, but the only thing that confuses me is the minus on the answer

hero (hero):

You can read more about it here: http://www.regentsprep.org/Regents/math/algtrig/ATT7/phaseshift.htm

hero (hero):

The general form of sinusoidal function is actually A cos(B(x - C)) + D Kinda confusing, but you factor out 2 then write it like this: 2(x - (-Π/4))

hero (hero):

So the Phase Shift really is -Π/4

hero (hero):

I know what I told you earlier but 2(x + Π/4) = 2(x - (-Π/4)) I probably just confused you more

hero (hero):

Writing 2(x + Π/4) as 2(x - (-Π/4)) puts it in the general form of sinusoidal function.

OpenStudy (christos):

Everything you said occurs for a function of cos as well?

OpenStudy (christos):

@Hero

hero (hero):

I don't quite understand your question. The general form of sinusoidal function is Acos(B(x - C)) + D The methods I used to find Amplitude and Phase Shift will be the same for most sinusoidal functions depending on what form they are in. You have to be able to recognize if they are in general form or not.

OpenStudy (christos):

Thank you very much for you help